QUANTITATIVE MODELING OF THE ν_{3} INFRARED SPECTRUM OF SF_{6} IN AN Ar MATRIX

\underline{TAO PENG} and ROBERT J. LE ROY, Guelph-Waterloo Centre for Graduate Work in Chemistry and Biochemistry, University of Waterloo, Waterloo, Ontario N2L 3G1, Canada.

An infrared active polyatomic molecule has several vibrational modes, each of which has a characteristic frequency. If the molecule is trapped in a matrix of perturbing atoms, those vibrational frequencies will shift, and if the vibrational mode is degenerate, the perturbation may lift the degeneracy. Such shifts and splittings are due to the dependence of the chromophore/matrix-atom interaction potential on the internal vibrational motion of the chromophore. Applying a previously-developed model for the shifting and splitting of the triply degenerate ν_{3} mode of SF_{6} perturbed by a rare gas atom,^{a} we use Monte Carlo simulations to find the most stable configurations for SF_{6} in an Ar matrix and to predict the associated thermally averaged perturbed IR spectra. Since the experimental spectrum has 10 peaks^{b} while the triply degenerate ν_{3} mode of SF_{6} in a particular environment could have at most 3 peaks, the observed spectrum must be a combination of spectra for SF_{6} trapped in different types of lattice sites. A fit to experiment of simulated spectra generated from a family of lattice sites determines the relative importance of the various sites, and accurately reproduces the experimental spectrum.

a D. Eichenauer and R.J. Le Roy, J. Chem. Phys. 88, 2898 (1988).

b B.I. Swanson and L.H. Jones, J. Chem. Phys. 74, 3205 (1981).